OpenStudy (muzzack):

Part A: Use the properties of exponents to explain why 64 raised to the power of 1 over 2 is called the square root of 64. Part B: The length of a rectangle is 7 units and its width is square root of 7 unit. Is the area of the rectangle rational or irrational? Justify your answer.

OpenStudy (muzzack):

@AccessDenied @azolotor @asatt32 @asher1027 @Ashy98 @animefreakdesu @bbcream14 @whpalmer4 @timo86m @thomaster @Ellia201 @theEric @Yvlotde900 @iPwnBunnies @NaomiBell1997 @walkingdeadxx

OpenStudy (anonymous):

what are you suposed to do?

OpenStudy (muzzack):

well i dont get it why is it called the square root of 64

OpenStudy (muzzack):

for Part A

OpenStudy (anonymous):

ummm hmmmm i dont no this one im so sorry tho that i could not help

OpenStudy (muzzack):

could you fan me

OpenStudy (muzzack):

@whpalmer4 @wagnermar @warmachine68 @walkingdeadxx @Whitemonsterbunny17 @esshotwired @Evictu_FB @Ellia201 @Emllyk @escapeshipshape @eifhbiew @escapeshipshape @freakenpurplepizza @FaithSmith5 @fishfry @Faraaz @gianna @GabrielleL. @girl_things @Hero

OpenStudy (anonymous):

o.O

OpenStudy (muzzack):

i know the rational exponent has to do with it for Part A but i still dont get it

OpenStudy (anonymous):

SCREW YOU MATH AAAAAAAAAAAHHHHHHHHHHHHHHHHHH

OpenStudy (muzzack):

YES IKR

OpenStudy (muzzack):

UHHHHHHHH

OpenStudy (muzzack):

I sound so nerdy

OpenStudy (muzzack):

so any advice?

OpenStudy (anonymous):

idk :'(

OpenStudy (muzzack):

man like once they read my question POOF everyone left

OpenStudy (muzzack):

its ok

OpenStudy (anonymous):

o.o i hate meth ill give u a ink 2 answer

OpenStudy (anonymous):

hmmmm @TheOcean @Whitemonsterbunny17 @ElvenBlackout

OpenStudy (muzzack):

thx @bbcream14 and @Emllyk

OpenStudy (anonymous):

your welcome :)

OpenStudy (muzzack):

how about Part B

OpenStudy (muzzack):

and @wagnermar thx for staying and trying to understand

OpenStudy (muzzack):

its hard for everyone

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

no problemo :)

OpenStudy (muzzack):

you guys are so awesome i fanned all of you @Emllyk @wagnermar @Whitemonsterbunny17 @bbcream14

whitemonsterbunny17 (whitemonsterbunny17):

Lol thanx!! You're awesome tooooo!!!!! :D

whitemonsterbunny17 (whitemonsterbunny17):

But I suck at math, so I can't really help much on this question lol :p

OpenStudy (muzzack):

its ok

OpenStudy (muzzack):

anyways i can always ask you for another thing

whitemonsterbunny17 (whitemonsterbunny17):

Yes, ofc!! (:

OpenStudy (muzzack):

ok let me close his and give you another one simpler

whitemonsterbunny17 (whitemonsterbunny17):

lol okeh :p

OpenStudy (whpalmer4):

"Use the properties of exponents to explain why 64 raised to the power of 1 over 2 is called the square root of 64. " \[64^{1/2} = \sqrt{64}\]The reason for this is simple. When we multiply exponential expressions with the same base, we keep the base and add the exponents: \[u^n*u^m = u^{n+m}\] Now, the square root of a number \(x\) can be defined as \[a = \sqrt{x}\]where \(a*a = x\) That means the square root is a number which multiplied by itself gives you the number of which you took the square root. Let's take that definition and square both sides of it (multiply each side by itself): \[a*a = \sqrt{x}*\sqrt{x}\]or \[a^2 = \sqrt{x}*\sqrt{x}\] Now, we can also write \(x = x^1\) by the properties of exponents, as every number without an exponent is assumed to have an exponent of 1. If \(a\) is the square root of \(x\), and \(a^2=x^1\), that means that to multiply those square roots together in exponential form (where the base is \(x\), the exponents must be such that they are equal and add to 1. The only way that is possible is if the exponent is \(1/2\), or \[\sqrt{x} = x^{1/2}\]Then when we multiply them together as we square both sides of \[a=x^{1/2}\]we get\[a^1*a^1 = x^{1/2}*x^{1/2}\]\[a^{1+1} = x^{1/2+1/2}\]\[a^2 = x^1\] Squaring as a term comes from what you do when you find the area of a square: you multiply the length by the width, and both are identical.